# biblio.bib

### v1

@@ 4 Discussion
After optimizing the synthesis, we show that the formulation of the emulsion can be optimized to combine different properties. First, the use of soybean oil and phospholipids make these soft matter tools inherently biocompatible. Second, the use of Tween 20 as a co-surfactant efficiently lowers the surface tension of the droplets, making them soft enough for stress sensing applications at cellular scales in a range of biological tissues, as evidenced by our injections inside organoids. These experiments ultimately allowed us to quantify a local stress of a few hundreds of pascals upon external deformation of the organoid. Third, this formulation still ensures the progressive destabilization of the inverted emulsion, which allows for self-functionalization and drug delivery properties of the droplets. The destabilization is not necessarily complete, as shown by the remaining inner water droplets visible in the injected droplet in the zebrafish embryos. However, using fluorescent binders or molecules still allows one to track precisely the extent of surface functionalization and the repartition of the binders in a given tissue. Moreover, these remaining inner droplets may be turned into an advantage. Indeed, the presence of discrete patches in the droplets or on their surface can be used as a referential for posture measurements. This could for instance allow to track events such as droplet rotations inside biological tissues.
In addition to this, the self-functionalization of the surface through phospholipids offers a wide platform for functionalization. Here we used biotinylated lipids that can be decorated with streptavidin and further grafted with any biotinylated molecule. One can also use phospholipids carrying a NTA-Ni group allow for the grafting of poly-histidine tagged molecules [Pontani2016a] or even phospholipids that offer a click chemistry platform on their hydrophilic head.
This simple, yet versatile, soft matter tool could therefore be used for a range of applications inside biological materials. Importantly, these will allow to perturb the biological systems, through specific binders presentation or local drug release, and directly measure the resulting effects of these perturbations on local stresses.
@@ Acknowledgements
The authors thank Marco Ribezzi and Victor Zahoransky for their help with the use of the SPG membrane connectors. Marie Anne Breau and Lea-Laetitia Pontani acknowledge financial support from Agence Nationale de la Recherche (MECAMATRIX, ANR-23-CE13-0025). Olga Vasiljevic aknowledges financial support from the i-Bio international PhD Program.

### v3

@@ 4 Discussion
After optimizing the synthesis, we show that the formulation of the emulsion can be tuned to combine different properties. First, the use of soybean oil and phospholipids make these soft matter tools inherently biocompatible. Second, the use of Tween 20 as a co-surfactant efficiently lowers the surface tension of the oil droplets, making them soft enough for stress sensing applications at cellular scales in a range of biological tissues, as evidenced by our injections inside organoids. These experiments ultimately allowed us to quantify a local stress of a few hundreds of pascals upon external deformation of the organoid. Third, this formulation still ensures the progressive destabilization of the inverted emulsion, which allows for self-functionalization and drug delivery properties of the sensors. The destabilization is not necessarily complete, as shown by the remaining inner water droplets visible in the injected oil droplet in the zebrafish embryos. However, using fluorescent binders or molecules still allows one to track precisely the extent of surface functionalization and the repartition of the binders in a given tissue. Moreover, these remaining inner water droplets may be turned into an advantage. Indeed, the presence of discrete patches in the oil droplets or on their surface can be used as a referential for posture measurements. This could for instance allow to track events such as rotations inside biological tissues.
In addition to this, the self-functionalization of the surface through phospholipids offers a wide platform for functionalization. Here we used biotinylated lipids that can be decorated with streptavidin and further grafted with any biotinylated molecule. One can also use phospholipids carrying a NTA-Ni group allow for the grafting of poly-histidine tagged molecules [51] or even phospholipids that offer a click chemistry platform on their hydrophilic head.
Importantly, these sensors could allow to decipher biological processes through different approaches. First, comparing sensors that interact with their surroundings through adhesive or passive contacts could help discriminate between different stress sources. For instance, whether a tissue is locally compressed laterally or pulled axially can lead to the same extension. Adhesive sensors could not only measure the stresses associated to extension but inform on its origin. Additionally, the self-functionalizing sensors can be used through a perturbative approach. The release of local drugs of cellular activity could thus unravel the contribution of cell generated forces into specific processes. Finally, thanks to the general framework that we introduced regarding their mechanical characterization, one can rationally measure either local stresses applied on the sensors, or unravel global stresses applied on a material whose properties are know. Therefore, this simple, yet versatile, soft matter tool could therefore be used for a range of applications inside biological materials.
@@ Acknowledgements
The authors thank Marco Ribezzi and Victor Zahoransky for their help with the use of the SPG membrane connectors, Jerome Fresnais for his help with surface tension measurements, and Loren Jorgensen for her help with the agarose gel rheometry. The authors also thank France Lam and Chloe Chaumeton for their support at the IBPS imaging facility. Finally, the authors thank Raphael Voituriez for fruitful discussions regarding mechanical modelling. Marie Anne Breau and Lea-Laetitia Pontani acknowledge financial support from Agence Nationale de la Recherche (MECAMATRIX, ANR-23-CE13-0025). Olga Vasiljevic aknowledges financial support from the i-Bio international PhD Program.
@@
Data availability Data will be available in a public repository.
Author contributions LLP, MB and JF conceived the study. OV, NH and LLP designed the experiments. OV, NH, EW, AH, LK and CH performed experiments. OV and NH analyzed data. OV, NH, VB, JF, EW, MB and LLP interpreted the results. OV, NH and LLP wrote the manuscript with feedback from all authors. LLP supervised the project.